The industrial production of nitric acid (HNO3) is a multistep process. The first step is the oxidation of ammonia (NH3) over a catalyst with excess oxygen (O2) to produce nitrogen monoxide (NO) gas as shown by the unbalanced equation given here:

?NH3(g)+?O2(g)→?NO(g)+?H2O(g)

What volume of O2 at 912 mmHg and 41 ∘C∘C is required to synthesize 24.5 mol of NO?

750 mmHg = 1 bar

R = 0.08314 L bar mol-1 K-1

Express your answer numerically in litres.
Answer #1

4 NH3  + 5 O2 4 NO + 6 H2O

Number of moles of Oxygen needed = 5/4 × 24.5 mol = 30.625 mol

PV = n RT

912/750 bar × V = 30.625mol × 0.08314L.bar/K.mol × 314.15K

V = 657.8 L

Volume of O2 required = 658 L (Answer)